OpenStudy (crashonce):

If p=1/3(x+x^-1) and q=1/3(x-x^-1), then what is p^2 - pq

OpenStudy (crashonce):

\[\frac{ 1 }{3 }(x+x ^{-1} = p \]

OpenStudy (crashonce):

similarly for q

OpenStudy (crashonce):

@ganeshie8

OpenStudy (anonymous):

mult out and see... p and q are conjugates

OpenStudy (anonymous):

i'm thinking 2/9

OpenStudy (anonymous):

\[p ^{2}-pq=\left[ \frac{ 1 }{ 3 }\left( x+\frac{ 1 }{ x } \right) \right]^{2}-\left[ \frac{ 1 }{ 3 }\left( x+\frac{ 1 }{ x } \right) \right]\left[ \frac{ 1 }{ 3 }\left( x-\frac{ 1 }{ x } \right) \right]\]

OpenStudy (anonymous):

\(x\ne 0\)

OpenStudy (anonymous):

\[1-q/p=1-\frac{ \left( x+\frac{ 1 }{ x } \right) }{ \left( x-\frac{ 1 }{ x } \right) }\]

OpenStudy (anonymous):

\[q=\frac{ x+\frac{ 1 }{ x } }{ x-\frac{ 1 }{ x } }p-2p\]

OpenStudy (anonymous):

reverse plus and minus signs of p and q in last and second to last equations.

OpenStudy (crashonce):

but that doesn't help find the value of p^2-pq does it?

OpenStudy (crashonce):

@hartnn can u explain

hartnn (hartnn):

whats \((x+x^{-1})^2 = ... ?\)

OpenStudy (crashonce):

x^2 + 2 +1/x^2

hartnn (hartnn):

and \((x+x^{-1})(x-x^{-1})=.. ?\)

OpenStudy (crashonce):

x^2 + a/x^2

hartnn (hartnn):

you mean x^2 -1/x^2 ?

OpenStudy (crashonce):

yes

hartnn (hartnn):

\(\Large p^2-pq = (1/9)(x+x^{-1})^2 -(1/9)(x+x^{-1})(x-x^{-1})\) plug in values ...

hartnn (hartnn):

did you get that bdw ?

OpenStudy (crashonce):

uh how should I start to simplify

OpenStudy (crashonce):

@hartnn

hartnn (hartnn):

i just asked whether u got it ?

OpenStudy (crashonce):

not the answer but yes I understand

hartnn (hartnn):

you already calculated the terms \(p^2-pq = (1/9) [(x^2+2+x^{-2}) - (x^2-x^{-2})]\)

OpenStudy (crashonce):

oh yea

hartnn (hartnn):

simplify that

OpenStudy (crashonce):

is it 2/9 (x^-2 -1)

hartnn (hartnn):

no

hartnn (hartnn):

\(p^2-pq = (1/9) [(x^2+2+x^{-2}) - (x^2-x^{-2})] \\ p^2-pq = (1/9) [2+x^{-2} +x^{-2})] = \huge 2/9 (1+x^{-2})\)

OpenStudy (crashonce):

oops typo

hartnn (hartnn):

ok :) np